# Shahriar Noroozizadeh

### v1

@@ F.2 Empirical intuition of the dynamics
Crucially, we find that this can happen (a) even without a constraint on the dimensionality $m$ and (b) this requires no early-stopping (see Remark˜2 for a more nuanced discussion of this).
@@
Next, due to the negative eigenvalues of ${\bf{C}}(t)$, the embeddings $\bf{V}$} along the lowermost eigendirections quickly diminish, achieving our first condition. Note that this means we do not want early-stopping; unlike in the quadratic loss formulation of Karkada et al. [75], it is longer training that filters out the lower eigenvectors. (Although, the existence of a degenerate eigenvector complicates this; see Remark˜2). This achieves the first condition in Section˜F.2.
@@ F.3 Mathematical description
At time $t=0$, the embeddings ${\bf{V}}(0)$ are all random, and hence nearly orthogonal to each other i.e., ${\bf{V}}(0){\bf{V}}(0)^{T}\approx c{\bf{I}}$, where $c$ is some constant that depends upon the magnitude of the random initialization. Since ${\bf{P}}=\tt{row\_softmax}({\bf{V}}{\bf{V}}(t))$, for a sufficiently large $c$, ${\bf{P}}(0)\approx{\bf{I}}$, proving our claim. ∎
@@
From the above observation, we can conclude that the eigenvectors of the system match the Laplacian throughout training. This follows by how the updates reduce to muplications between matrices sharing the same eigenspaces.
@@
Now, we can establish the conditions in Section˜F.2, namely, the convergence of the null space of the co-efficient matrix from Section˜2.1.1, and then the convergence of the embedding vectors.
@@
Recall that ${\bf{C}}(t)={\bf{D}}^{-1}{\bf{A}}+({\bf{D}}^{-1}{\bf{A}})^{T}-({\bf{P}}+{\bf{P}}^{T})$. The eigenvalues of the first term lie approximately in [2,2][-2,2] from Proposition˜8, while that of the probability term lie in $[0,2]$, from Proposition˜7. Note that eigenvalues of the probability matrix all begin uniformly at $2$ in the beginning (as ${\bf{P}}(0)={\bf{I}}$, by ˜1 under ˜1), as a result of which the initial eigenvalues of ${\bf{C}}(t)$ start at or below 0.
While the embeddings are initialized orthogonally under ˜1, they become less orthogonal during training, leading to a gradual decrease of the diagonal self-probability terms in ${\bf{P}}+{\bf{P}}^{T}$. Intuitively, this also means that the eigenvalues of ${\bf{P}}+{\bf{P}}^{T}$ must themselves all decrease from the initial value of $2$ (based on Proposition˜7). In turn, the eigenvalues of ${\bf{C}}(t)$, which begin negative must gradually inch toward zero. The topmost eigenvalues—which are closest to zero—are the first to reach zero. [^7] Note that this is not straightforward to show. It is possible that the even if the initial eigenvalues are very close to zero, they approach 0 slower than farther off values.
@@
Given that ${\bf{C}}(t)$ have time-invariant eigenvectors (by Proposition˜6), this can be further simplified as
@@
Given that the eigenvalues in $\bf{\Lambda}$} are all less than or equal to zero (by Proposition˜9), the term $(1+\eta{\bf{\Lambda}}(t))$ must consist of a diagonal of values in $[0,1]$ for an appropriately small learning rate. Furthermore, as $T$ becomes large, the values of the top eigenvectors (which have the least eigenvalues, and therefore, the largest value of $1+\eta\lambda_{i}(t)$) must come to dominate. Then, as $t$ increases, we can express the embedding dimension as an affine combination of some top $K$ eigenvectors (where the coefficients depend on how the embedding dimension was initialized):

### v2

@@ F.2 Empirical intuition of the dynamics
Crucially, we find that this can happen (a) even without a constraint on the dimensionality $m$ and (b) this requires no early-stopping (see Remark˜5 for a more nuanced discussion of this).
@@
Next, due to the negative eigenvalues of ${\bf{C}}(t)$, the embeddings $\bf{V}$} along the lowermost eigendirections quickly diminish, achieving our first condition. Note that this means we do not want early-stopping; unlike in the quadratic loss formulation of Karkada et al. [83], it is longer training that filters out the lower eigenvectors. (Although, the existence of a degenerate eigenvector complicates this; see Remark˜5). This achieves the first condition in Section˜F.2.
@@ F.3 Mathematical description
At time $t=0$, the embeddings ${\bf{V}}(0)$ are all random, and hence nearly orthogonal to each other i.e., ${\bf{V}}(0){\bf{V}}(0)^{T}\approx c{\bf{I}}$, where $c$ is some constant that depends upon the magnitude of the random initialization. Since ${\bf{P}}=\tt{row\_softmax}({\bf{V}}{\bf{V}}^{T}(t))$, for a sufficiently large $c$, ${\bf{P}}(0)\approx{\bf{I}}$, proving our claim. ∎
@@
From the above observation, we can conclude that the eigenvectors of the system match the Laplacian throughout training. This follows by how the updates reduce to multiplications between matrices sharing the same eigenspaces.
@@
Now, we can establish the conditions in Section˜F.2, namely, the convergence of the null space of the co-efficient matrix from Section˜2.2, and then the convergence of the embedding vectors.
@@
Recall that ${\bf{C}}(t)={\bf{D}}^{-1}{\bf{A}}+({\bf{D}}^{-1}{\bf{A}})^{T}-({\bf{P}}+{\bf{P}}^{T})$. The eigenvalues of the first term lie approximately in [2,2][-2,2] from Proposition˜9, while that of the probability term lie in $[0,2]$, from Proposition˜8. Note that eigenvalues of the probability matrix all begin uniformly at $2$ in the beginning (as ${\bf{P}}(0)={\bf{I}}$, by ˜1 under ˜1), as a result of which the initial eigenvalues of ${\bf{C}}(t)$ start at or below 0.
While the embeddings are initialized orthogonally under ˜1, they become less orthogonal during training, leading to a gradual decrease of the diagonal self-probability terms in ${\bf{P}}+{\bf{P}}^{T}$. Intuitively, this also means that the eigenvalues of ${\bf{P}}+{\bf{P}}^{T}$ must themselves all decrease from the initial value of $2$ (based on Proposition˜8). In turn, the eigenvalues of ${\bf{C}}(t)$, which begin negative must gradually inch toward zero. The topmost eigenvalues—which are closest to zero—are the first to reach zero. [^15] Note that this is not straightforward to show. It is possible that even if the initial eigenvalues are very close to zero, they approach 0 slower than farther off values.
@@
Given that ${\bf{C}}(t)$ have time-invariant eigenvectors (by Proposition˜7), this can be further simplified as
@@
Given that the eigenvalues in $\bf{\Lambda}$} are all less than or equal to zero (by Proposition˜10), the term $(1+\eta{\bf{\Lambda}}(t))$ must consist of a diagonal of values in $[0,1]$ for an appropriately small learning rate. Furthermore, as $T$ becomes large, the values of the top eigenvectors (which have the least eigenvalues, and therefore, the largest value of $1+\eta\lambda_{i}(t)$) must come to dominate. Then, as $t$ increases, we can express the embedding dimension as an affine combination of some top $K$ eigenvectors (where the coefficients depend on how the embedding dimension was initialized):